# How to Upscale Neural Networks with Scaling Law? A Survey and Practical Guidelines

### v1

@@ 1 Introduction
While prior surveys have made valuable contributions to understanding scaling laws, they have primarily focused on specific aspects of the scaling phenomenon (See Table 1). Choshen et al. (2024) emphasized statistical best practices for estimating and interpreting scaling laws using training data, while Li et al. (2024b) emphasized on methodological inconsistencies in scaling studies and the reproduction crisis in scaling laws. Our survey distinguishes itself by offering comprehensive coverage of architectural considerations, data scaling implications, and inference scaling – areas that previous surveys either overlooked or addressed only partially.
@@ 2 Taxonomy of neural scaling laws
with appropriate scaling parameters and fitting parameters for different scaling parameter . Figure 3b highlights that the number of model parameters and data size are the most common used scaling factors. The exact forms of all the scaling laws are highlighted in Table 7 of Appendix B. Among all the tasks, Figure 3c suggests that language generation is the most common task used for developing these scaling laws, where the training cross-entropy loss is widely used to fit the laws. Based on the values obtained empirically, the scaling laws are fitted with non-linear optimization, most commonly by running algorithms like least square and BFGS (Broyden-Fletcher-Goldfarb-Shanno). Statistical methods like goodness-of-fit metrics are used to validate the correctness of the fitted curves. We elaborate on the evaluation of neural scaling laws in Appendix A.
In the following sections, we review the existing literature on neural scaling across various domains.
2.1 Scaling laws of language models
| Modality | Paper | Key insights | Applicability |
[…]
2.2 Scaling laws in other modalities
2.3 Scaling laws for domain adaptation
Fine-tuning strategies scale differently depending on dataset size. Parameter-efficient fine-tuning (PEFT) is well-suited for small datasets, low-rank adaptation (LoRA) (Hu et al., 2021) performs best for mid-sized datasets, and full fine-tuning is most effective for large datasets. However, PEFT methods provide better generalization in large models, making them attractive alternatives to full-scale fine-tuning (Zhang et al., 2024).
| Paper | Key insights | Applicability |
[…]
2.4 Scaling laws for model inference
| Paper | Key insights | Applicability |
[…]
2.5 Scaling laws for efficient models
2.6 Data scaling laws
| Paper | Key insights | Applicability |
[…]
3 Practical guidelines for utilizing scaling laws
Large-scale adoption of neural models requires careful determination of several key factors:
Model size vs. data size: How should the number of parameters and dataset size be balanced to achieve optimal performance?
Compute budget: Given a fixed computational budget, how can training and inference be optimized for efficiency?
Training strategy: Should longer training duration be prioritized over increased model size, especially in data-constrained environments?
Architecture choice: Should the model be dense or sparse (e.g., MoE) to optimize performance for a given budget?
Inference efficiency: What test-time strategies, such as retrieval augmentation or tree search, can be employed to enhance model output quality?
Quantization and pruning: How can quantization and structured sparsity be used to minimize memory footprint while maintaining accuracy?
Based on the selections of the above choices, one can follow the steps highlighted in Figure 4 for developing large-scale neural models. Figure 4a illustrates that for developing large-scale models with multimodal capabilities, multimodal scaling laws Aghajanyan et al. (2023) can be followed. For developing domain-specific nuanced application, one can follow D-CPT law Que et al. (2024), which provides a comprehensive guideline for domain-continuous pre-training. For developing general-purpose LLMs, one may need to follow Kaplan/Chinchilla scaling laws, if sufficient data is available and the model choice is dense; otherwise, MoE scaling laws Krajewski et al. (2024) be utilized to develop sparser models. If sufficient data is not available, one can either train the model for longer epochs, following Sardana et al. (2024), or use data mixing strategies Ye et al. (2024) for deriving optimal data composition for pre-training the model.
Post-training, inference strategies (c.f. Figure 4b) can be followed to ensure that the model is utilized efficiently for the downstream applications. If the high memory consumption of the model is a concern, compression strategies like pruning Chen et al. (2024b) or low-precision quantization Cao et al. (2024) can be used. To make larger models inference efficient, inference scaling laws Wu et al. (2024) or test-time scaling laws Chen et al. (2024a) can be utilized. Appendix D elaborates more practical guidelines in terms of using the existing literature of neural scaling laws for developing large-scale neural models.
@@ 4 Criticisms of scaling laws
Despite their significant impact, many studies on scaling laws suffer from limited reproducibility (see Table 8 in Appendix C) due to proprietary datasets, undisclosed hyperparameters, and undocumented training methodologies. The inability to replicate results across different computing environments raises concerns about their robustness. Large-scale experiments conducted by industry labs often depend on private infrastructure, making independent verification challenging. This lack of transparency undermines the reliability of scaling law claims and highlights the urgent need for open benchmarks and standardized evaluation frameworks to ensure reproducibility.
5 Future recommendations
@@
Inference-aware scaling: Scaling laws should incorporate test-time computation, as compute-efficient inference strategies (e.g., iterative refinement, tree search, retrieval-based augmentation) can allow smaller models to outperform larger ones. Future research should focus on balancing training vs. inference compute costs to develop models that scale efficiently in real-world applications.
Compute-optimal model selection: Scaling laws should not only predict performance improvements but also guide model selection given a fixed compute budget. Future work should explore multi-objective optimization frameworks that balance performance, energy efficiency, and cost to drive more sustainable AI development.
Efficient data scaling and pruning: The optimization of model scaling necessitates a shift from volume-based to quality-focused data selection. Future frameworks should prioritize informative examples and integrate diversity metrics to enhance generalization, moving beyond simple dataset expansion.
@@ 6 Conclusion
This survey provided a comprehensive analysis of neural scaling laws, exploring their theoretical foundations, empirical findings, and practical implications. It synthesized insights across various modalities, including language, vision, multimodal learning, and reinforcement learning, to uncover common trends and deviations from traditional power-law scaling. While early research established predictable relationships between model size, dataset volume, and computational resources, more recent studies have shown that these relationships are not universally applicable. Sparse architectures, retrieval-augmented models, and domain-specific adaptations often exhibit distinct scaling behaviors, challenging the notion of uniform scalability. Furthermore, advancements in fine-tuning, data pruning, and efficient inference strategies have introduced new perspectives on compute-optimal scaling. Despite their significance, scaling laws remain an evolving area of research, requiring further refinement to address real-world deployment challenges and architectural innovations.
@@ A.5 Limitations of fitting techniques
Appendix B Other scaling laws
B.1 Scaling laws for reinforcement learning
@@
B.2 Sparse autoencoders
B.3 Graph neural networks
Unlike in computer vision and natural language processing, where larger datasets typically improve model generalization, graph self-supervised learning (Graph SSL) methods fail to exhibit expected scaling behavior and performance fluctuates unpredictably across different data scales (Ma et al., 2024). However, self-supervised learning pretraining loss does scale with more training data, but this improvement does not translate to better downstream performance. The scaling behavior is method-specific, with some approaches like InfoGraph showing more stable scaling than others like GraphCL.
@@ Appendix C Reproducibility of scaling laws papers
Appendix D Additional research questions and guidelines
RQ1. How does the relationship between model size, training data, and compute efficiency influence LLM performance across different tasks?
For optimal LLM scaling, Kaplan et al. (2020) proposed that model performance follows a power law:
where is the loss, is the model size, is the dataset size. They suggest an increase of 8 model size requires a 5 data increase to maintain efficiency:
Hoffmann et al. (2022) refined this by adding an irreducible loss term:
They also suggested balancing model and data scaling equally:
Small LMs require a higher data-to-model ratio () and benefit from continuous training and domain adaptation (Hu et al., 2024). Broken Neural Scaling Laws (BNSL) describe non-monotonic scaling behaviors, modeled by a piecewise function (Caballero et al., 2023):
where represents the transition threshold. When data is constrained, smaller models with more epochs perform better, with performance gains plateauing around 4 epochs but still beneficial up to 16 epochs (Sardana et al., 2024; Muennighoff et al., 2023):
where is the number of epochs and is the saturation point. To maximize efficiency, prioritizing data-to-parameter ratios over raw model expansion ensures that smaller, well-trained models outperform larger undertrained ones. High-quality, diverse data and efficient training remain keys to scalable and robust AI models.
RQ2. How do different model architectures scale differently, and what are the architectural bottlenecks?
According to Tay et al. (2022), the vanilla Transformer consistently demonstrates superior scaling properties compared to other architectures, even though alternative designs might perform better at specific sizes. Architectural bottlenecks manifest differently across these designs. For instance, linear attention models like Performer and lightweight convolutions show inconsistent scaling behavior, while ALBERT demonstrates negative scaling trends. This finding helps explain why most LLMs maintain relatively standard architectures rather than adopting more exotic variants.
where is the performance metric, represents compute, and are fitting parameters. Furthermore, Zhai et al. (2022) revealed that ViT reveals that these models exhibit double saturation –performance plateaus at both very low and very high compute levels, suggesting architectural limitations specific to the vision domain (Equation 9). However, as shown by Li et al. (2024a), simply scaling up vision encoders in multimodal models does not consistently improve performance, indicating that architectural scaling benefits are not uniform across modalities.
These findings suggest several practical implications for practitioners. First, when selecting an architecture for scaling, it is crucial to evaluate its scaling behavior rather than just its performance at a single size point. The indicates that practitioners should be cautious about adopting architectures that drastically modify core Transformer components, especially for large-scale deployments. Additionally, the optimal architecture may depend on the specific compute scale, what works best at a small scale may not be ideal for large models.
RQ3. What are the optimal data mixture strategies to achieve maximum performance at different scales?
Recently, techniques like REGMIX (Liu et al., 2024) and AUTOSCALE (Kang et al., 2024) have been proposed to optimize data mixtures for various model scales. Ye et al. (2024) introduced the following predictive model (Equation 10), demonstrating that performance can be estimated using exponential functions that combine proportions from different domains, enabling pre-optimization of training mixtures before full-scale training begins.
where denotes validation loss for domain , represents the proportions of domains, and , and are fitting parameters.
Additionally, domain interactions in continual pre-training have proven to be more complex than previously assumed. The optimal mixture ratio between general and domain-specific data follows predictable scaling patterns, as described by the domain-continual pre-training law (Que et al., 2024) in Equation 11.
where represents the number of model parameters, is the dataset size, is the mixture ratio, and are fitting parameters.
For practitioners, research suggests adopting a dynamic approach to data mixture optimization that adjusts composition with scale rather than maintaining fixed ratios throughout training. Carefully considering compute budgets and target domains when designing training strategies can significantly improve model performance while reducing computational costs through more efficient data utilization.
RQ4. How does test-time compute scaling compare to model size scaling in terms of efficiency and performance?
Recent research examining the relationship between test-time computation and model size scaling has revealed key insights. Brown et al. (2024) proposed that repeated sampling during inference significantly enhances model performance, with coverage (fraction of problems solved) following an exponentiated power law relationship with the number of samples , as described by Equation 12.
where are fitting parameters.
Further exploration by Wu et al. (2024) suggested that employing sophisticated test-time computation strategies (such as iterative refinement or tree search) with smaller models may be more cost-effective than using larger models with simple inference methods. Their work establishes a relationship between inference computational budget and optimal model size for compute-efficient inference, expressed as:
These findings indicate that hybrid approaches combining model size optimization with test-time computation strategies may offer the best trade-off for organizations deploying language models. Practitioners should assess the complexity of their target tasks and consider adaptive computation strategies that adjust inference resources based on problem difficulty. This approach can lead to significant cost savings while maintaining performance, particularly for routine tasks that do not require the full capacity of larger models.
RQ5. How does scaling fine-tuning parameters affect performance on downstream tasks?
Hernandez et al. (2021) proposed scaling laws for transfer by fine-tuning decoder-only transformer models on python code. They introduced a concept of effective data transferred , i.e., the amount of additional python data that a model of the same size trained on only python would have needed to achieve the same loss on python as a model pre-trained on language, as a function of fine tuning data . The law is given as :
Along the same tracks, Lin et al. (2024a) proposed a rectified scaling law given by Equation 15. They introduced the term pre-learned data size that indicates how much amount of downstream data a model has learned from pre-training:
where is the fine-tuning data size and are fitting parameters.
Abnar et al. (2021) predicted, downstream error for image recognitions tasks on ViTs and ResNets as a function of upstream error , given by the equation:
This was further explored by Mikami et al. (2021), modelling pre-training data size, consisting of syntheic dataset to predict downstream error with following equation:
Chen et al. (2024c) proposed a framework for predicting downstream performance, called FLP method which first, maps FLOPS to training loss and then maps the pre-training loss to downstream task performance. Zhang et al. (2024) studied scaling effects under parameter-efficient fine-tuning (PEFT and LoRA). They found that LoRA and prompt tuning (Lester et al., 2021) scale poorly when PEFT parameters are increased, although LoRA has superior training stability. LLM fine-tuning scaling is task- and data-dependent, making the choice of the best fine-tuning approach for a downstream task non-trivial. They proposed a multiplicative joint scaling law:
where denotes fine-tuning data, denotes either LLM model size or pretraining data size or PEFT parameter size, and are fitting parameters.
RQ5. How do sparse model scaling laws (e.g., Mixture-of-Experts, Pruning) compare to dense model scaling laws in terms of efficiency and performance?
Recent research has revealed important insights into how sparse models scale compared to dense models. Frantar et al. (2023) demonstrated that sparsity acts as a multiplicative constant to model size scaling rather than fundamentally altering the scaling behavior (Equation 19). This implies that while sparse models can be more efficient, they follow similar underlying trends as dense models.
where represents the validation loss, represents sparsity, is the number of non-zero parameters, is the training dataset size, and are fitting parameters.
For Mixture-of-Experts (MoE) models, Clark et al. (2022) proposed the following law for a constant dataset size:
where is the final loss, represents the number of parameters an input encounters, and is the expansion rate, are fitting parameters. The scalability of MoE models is constrained by the need to balance training samples with computational resources for optimal utilization. As noted by Kaplan et al. (2020) and Hoffmann et al. (2022), maintaining a constant dataset size while increasing model size can lead to undertraining, limiting model performance.
To address this, Yun et al. (2024) extended Equation 20 to incorporate dataset size:
where are fitting parameters.
Furthermore, Krajewski et al. (2024) introduced a granularity parameter () to refine Equation 4 for MoE models. Their research demonstrates that well-configured MoE models, with optimal granularity and training parameters, are consistently more efficient than dense Transformers across all scales:
where are fitting parameters.
Additionally, Chen et al. (2024b) introduced the Law (Equation 23) for understanding the performance of pruned models post-training. Their findings suggest that post-training loss after pruning is influenced by four key factors: the pre-pruning model size, the number of post-training tokens, the pruning rate, and the model’s pre-pruning loss. This provides a structured approach for optimizing sparse model training.
where is the uncompressed model loss, is the pruning ratio, is the pre-pruning model size, represents the number of post-training tokens, and are fitting parameters.
These findings collectively suggest that properly implemented sparse architectures could provide a more compute-efficient pathway for scaling language models. This is particularly relevant as the field moves toward sustainable AI practices. The key to success lies in careful optimization – whether through precise expert size configuration in MoE models or strategic post-training approaches for pruned models.
RQ7. What are the optimal quantization strategies to maintain model performance while reducing inference cost?
According to Dettmers and Zettlemoyer (2023), 4-bit precision appears to be the optimal sweet spot for maximizing model performance while minimizing model size. Additionally, research on scaling with mixed quantization (Cao et al., 2024), demonstrated that larger models can handle higher quantization ratios while maintaining performance, following an exponential relationship where larger models require exponentially fewer high-precision components to maintain a given performance level.
Kumar et al. (2024) developed a unified scaling law (Equation 24) that predicts both training and post-training quantization effects. It further suggests that effects of quantizing weights, activations, and attention during training are independent and multiplicative.
These findings suggest opportunities for developing specialized architectures optimized for ultra-low precision at scale. The discovery that quantization effects are independent across model components points to potential benefits of dynamic precision adjustment based on computational needs. Combined with insights about the relationship between training precision and compute budgets, this could lead to more efficient training approaches and architectures inherently designed for robust quantization performance.
RQ8. Do scaling laws hold in multimodal models, and how does modality interaction affect performance at scale?
Henighan et al. (2020) first explored the scaling laws for multimodals, proposing the form :
where can be either model size, data or compute. This was further improved by Alabdulmohsin et al. (2022) by proposing the form :
While assumes that model performance follows a simple power law relationship (meaning performance improves at a constant rate as you add more training data), real models often don’t behave so cleanly. introduces additional flexibility through a sigmoid-like shape that can smoothly transition between different phases of learning.
Further research has revealed fascinating dynamics in how different modalities interact as models grow larger. According to Aghajanyan et al. (2023), there exists an inherent competition between modalities in smaller models, which can transform into synergy as model size increases after crossing “competition barrier”. They used Chinchilla form (Equation 4) as a uni-modal scaling law, and improvised on it to propose a bi-modal scaling law:
where denotes loss, N denotes model size, and denote two datasets being used and represent the maximum level of synergy.
The findings indicate that for multimodal models, practitioners should ensure sufficient model size to enable synergistic rather than competitive interaction between modalities, while emphasizing that scaling strategies should prioritize data quality and modality alignment over mere size increases. High-quality image-text pairs and thoughtful consideration of inter-modal relationships in architecture design are more crucial than simply scaling up individual components.

### v2

@@ 1 Introduction
While prior surveys have made valuable contributions to understanding scaling laws, they have primarily focused on specific aspects of the scaling phenomenon (See Table 1). Choshen et al. (2024) emphasized statistical best practices for estimating and interpreting scaling laws using training data, while Li et al. (2024b) emphasized on methodological inconsistencies and reproduction crisis in existing scaling laws. Our survey distinguishes itself by offering comprehensive coverage of architectural considerations, data scaling implications, and inference scaling – areas that previous surveys either overlooked or addressed only partially.
@@ 2 Taxonomy of neural scaling laws
with appropriate scaling parameters and fitting parameters for different scaling parameter . Figure 3b highlights that the number of model parameters and data size are the most common used scaling factors. The exact forms of all the scaling laws are highlighted in Table 10 of Appendix B. Among all the tasks, Figure 3c suggests that language generation is the most common task used for developing these scaling laws, where the training cross-entropy loss is widely used to fit the laws. Based on the values obtained empirically, the scaling laws are fitted with non-linear optimization, most commonly by running algorithms like least square and BFGS (Broyden-Fletcher-Goldfarb-Shanno). Statistical methods like goodness-of-fit metrics are used to validate the correctness of the fitted curves. We elaborate on the evaluation of neural scaling laws in Appendix A. In the following sections, we review the existing literature on neural scaling across various domains.
Model scaling includes both parameter and data scaling. Parameter scaling is often studied in decoder-only Transformers (Kaplan et al., 2020; Hoffmann et al., 2022), with newer works addressing small and efficient models (Hu et al., 2024; Clark et al., 2022). These studies establish power-law relationships between loss and model size or compute (Equation 1). In parallel, data scaling research has proposed laws for optimizing mixtures (Ye et al., 2024), repeated training exposures (Muennighoff et al., 2023), vocabulary size (Tao et al., 2024), and knowledge capacity (Allen-Zhu and Li, 2024).
Pre-training scaling laws extend beyond language to vision and multimodal settings. Vision models exhibit power-law scaling that saturates at large compute (Zhai et al., 2022), while multimodal models demonstrate competition-to-synergy transitions as scale increases (Aghajanyan et al., 2023).
Post-training scaling captures fine-tuning and transfer learning behaviors. Transfer scaling shows larger pre-trained models yield better generalization with limited downstream data (Hernandez et al., 2021). Recent works propose scaling laws for PEFT (Zhang et al., 2024), downstream loss prediction (Chen et al., 2024c), and early stopping (Lin et al., 2024a).
Inference scaling explores compute-efficient strategies during model deployment. Adaptive test-time compute (Chen et al., 2024a; Brown et al., 2024) and retrieval augmentation (Shao et al., 2024) allow small models to rival larger ones. Inference-specific scaling laws characterize the tradeoff between sampling cost and performance (Wu et al., 2024).
Efficient model scaling addresses sparsity, quantization, and distillation. Sparse and MoE models provide multiplicative efficiency gains (Krajewski et al., 2024), while pruning and quantization laws enable compute-aware compression (Chen et al., 2024b; Cao et al., 2024).
Scaling behavior in reinforcement learning (RL) diverges from language or vision tasks. In single-agent RL, performance scales sublinearly with model size and environment interaction (Hilton et al., 2023). Horizon length, rather than task difficulty, determines scaling efficiency. In multi-agent games, predictable scaling laws govern compute-to-performance relationships, but generalization to complex domains like Chess or Go remains limited (Neumann and Gros, 2023). Meanwhile, graph neural networks (GNNs) lack stable scaling laws; despite self-supervised loss improving with more data, downstream performance often fluctuates unpredictably (Ma et al., 2024).
Finally, the taxonomy captures two outer branches: commendations, such as practical data laws and compression-aware training (Liu et al., 2024), and criticisms, which question the generalizability and reproducibility of scaling laws (Sorscher et al., 2023; Diaz and Madaio, 2024). Detailed discussion on these scaling law studies are provided in Appendix B.
In the next section, we formulate key research questions (mapping between the taxonomy and research questions highlighted in Table 2) derived from these studies and present practical guidelines for leveraging scaling laws in real-world model development.
| Taxonomy Node | Addressed RQs |
| Model scaling | RQ1, RQ2, RQ8 |
| Data scaling | RQ3 |
| Post-training scaling | RQ5 |
| Inference scaling | RQ4 |
| Efficient and compressed model scaling | RQ6, RQ7 |
3 Research questions and guidelines
Grounded in the taxonomy of neural scaling laws (Figure 2), we identify key research questions spanning six dimensions: model scaling, architectural bottlenecks, inference scaling, data scaling, post-training strategies, and efficient model design. For each, we synthesize multiple studies to extract overarching patterns, identify conflicting evidence, and propose actionable guidelines for researchers and practitioners navigating large-scale model development.
RQ1. Importance on model and pre-training data size on performance. [taxonomy: model scaling pre-training]
Kaplan et al. (2020) established a power-law relationship:
Hoffmann et al. (2022) refined this into a compute-optimal formulation:
Recent research has challenged linear extrapolations. Muennighoff et al. (2023) and Sardana et al. (2024) showed that training small models longer can outperform larger models, especially under constrained data. Caballero et al. (2023) proposed Broken Neural Scaling Laws (BNSL):
RQ2. Scaling behaviors for different neural architectures. [taxonomy: model scaling pre-training architecture]
According to Tay et al. (2022), the vanilla Transformer consistently demonstrates superior scaling properties (, where is the performance metric, represents compute, and are fitting parameters) compared to other architectures, even though alternative designs might perform better at specific sizes. Architectural bottlenecks manifest differently across these designs. For instance, linear attention models like Performer and Lightweight Convolutions show inconsistent scaling behavior, while ALBERT demonstrates negative scaling trends. This finding helps explain why most LLMs maintain relatively standard architectures rather than adopting more exotic variants. Furthermore, Zhai et al. (2022) revealed that ViT reveals that these models exhibit double saturation, where performance plateaus at both very low and very high compute levels, suggesting architectural limitations specific to the vision domain (Equation 5). However, as shown by Li et al. (2024a), simply scaling up vision encoders in multimodal models does not consistently improve performance, indicating that architectural scaling benefits are not uniform across modalities.
RQ3. Data strategies for performance scaling. [taxonomy: data scaling]
Ye et al. (2024) proposed an exponential model for data mixing:
while Liu et al. (2024) and Kang et al. (2024) developed proxy models (REGMIX, AUTOSCALE) to pre-optimize mixtures. The Domain-Continual Pretraining (D-CPT) law (Que et al., 2024) provides a theoretical grounding on optimal mixture ratio between general and domain-specific data :
where represents the number of model parameters, is the dataset size, is the mixture ratio, are fitting parameters.
RQ4. Test-time scaling for better scaling efficiency. [taxonomy: model scaling inference scaling]
Recent research examining the relationship between test-time computation and model size scaling has revealed key insights. Brown et al. (2024) proposed that repeated sampling during inference significantly enhances model performance, with coverage (fraction of problems solved) following an exponentiated power law relationship with the number of samples , , where are fitting parameters. Further exploration by Wu et al. (2024) suggested that employing sophisticated test-time computation strategies (such as iterative refinement or tree search) with smaller models may be more cost-effective than using larger models with simple inference methods. Their work establishes a relationship between inference computational budget and optimal model size for compute-efficient inference, expressed as .
RQ5. Scaling behaviors of model fine-tuning. [taxonomy: model scaling post-training scaling]
Fine-tuning scaling reflects how pre-trained models adapt across tasks and domains. Hernandez et al. (2021) introduced a transfer scaling law based on effective data transferred :
while Lin et al. (2024a) refined this with a rectified law:
modeling diminishing returns from fine-tuning beyond a pre-learned threshold. In vision, Abnar et al. (2021) linked downstream error to upstream error:
and Mikami et al. (2021) connected downstream accuracy to synthetic pretraining data size:
FLOPS to Loss to Performance (FLP) method (Chen et al., 2024c) predicted downstream performance from pretraining FLOPs, and Zhang et al. (2024) showed LoRA scales nonlinearly under PEFT:
RQ6. Scaling efficiency and performance for sparse and efficient models. [taxonomy: model scaling model compression]
As the demand for resource-efficient models grows, sparse architectures such as pruned networks and MoEs have emerged as promising alternatives to dense Transformers. These models aim to preserve the performance benefits of scale while reducing compute and memory overhead. Frantar et al. (2023) proposed a general sparse scaling law showing that sparsity acts as a multiplicative efficiency factor rather than changing the fundamental scaling behavior:
where is sparsity, is the number of non-zero parameters, and is dataset size. In MoE models, where only a subset of parameters is activated per input, Clark et al. (2022) proposed a loss scaling relationship incorporating both model size and expert count:
with denoting the expansion factor. This formulation was extended by Yun et al. (2024) to include dataset size:
These results emphasize that scaling MoEs effectively requires balancing expert granularity with sufficient training data. Toward this, Krajewski et al. (2024) introduced a granularity parameter to refine the Chinchilla-style formulation:
In parallel, structured pruning approaches have been formalized through the law (Chen et al., 2024b), which relates post-pruning loss to pre-pruning model size , pruning ratio , and post-training token count :
where is the uncompressed model loss, is the pruning rate, is the pre-pruning model size, represents the number of post-training tokens, and are fitting parameters.
RQ7. Model scaling with low-precision quantization. [taxonomy: model scaling model compression quantization ]
According to Dettmers and Zettlemoyer (2023), 4-bit precision appears to be the optimal sweet spot for maximizing model performance while minimizing model size. Additionally, research on scaling with mixed quantization (Cao et al., 2024), demonstrated that larger models can handle higher quantization ratios while maintaining performance, following an exponential relationship where larger models require exponentially fewer high-precision components to maintain a given performance level. Kumar et al. (2024) developed a unified scaling law (Equation 18) that predicts both training and post-training quantization effects. It further suggests that effects of quantizing weights, activations, and attention during training are independent and multiplicative.
RQ8. Beyond modalities: scaling for multimodal models. [taxonomy: model scaling multimodal models]
Multimodal scaling behavior builds upon, but does not replicate, unimodal trends. Henighan et al. (2020) first proposed multimodal scaling using , where represents model size, data, or compute. Alabdulmohsin et al. (2022) refined this into a more flexible sigmoid-like form:
allowing transitions across saturation regimes. Aghajanyan et al. (2023) observed that smaller multimodal models exhibit competition between modalities, while larger models cross a “competition barrier” and become synergistic. They proposed a bimodal generalization of the Chinchilla law:
where captures the degree of positive interaction between modalities and .
While the research questions synthesized above highlight the strengths and practical applications of neural scaling laws, they also expose several limitations, especially in their generalizability, reliability under constraints and applicability to modern model designs. In the next section, we critically examine these limitations and discuss the foundational assumptions that may no longer hold as models evolve.
@@ 4 Criticisms of scaling laws
Despite their significant impact, many studies on scaling laws suffer from limited reproducibility (see Table 11 in Appendix C) due to proprietary datasets, undisclosed hyperparameters, and undocumented training methodologies. The inability to replicate results across different computing environments raises concerns about their robustness. Large-scale experiments conducted by industry labs often depend on private infrastructure, making independent verification challenging. This lack of transparency undermines the reliability of scaling law claims and highlights the urgent need for open benchmarks and standardized evaluation frameworks to ensure reproducibility. Furthermore, the field’s avoidance of rigorous scaling exponent analysis constitutes a critical oversight. While exponents indeed vary across models, datasets, and hyperparameters, this variability demands investigation rather than dismissal. This deliberate analytical gap undermines confidence in extrapolation claims and raises questions about whether observed scaling behaviors represent genuine properties or experimental artifacts.
5 Beyond Scale: Future Directions for Practical and Sustainable AI
@@
Reframing scaling laws for real-world constraints.
Future scaling laws must account for compute budgets, hardware latency, and energy consumption. This includes integrating training–inference trade-offs, evaluating real-world performance under quantization or pruning, and predicting effectiveness across resource-constrained environments.
Designing for downscaling.
Rather than building ever-larger models, the field should invest in scaling laws for small language models trained with optimal data, sparsity, and inference strategies. The emergence of 1-3B parameter models that rival 13B+ models (Hu et al., 2024) highlights the viability of compact yet performant systems.
Multi-objective scaling optimization.
Current scaling laws often predict accuracy at scale but ignore trade-offs between accuracy, compute, and robustness. Future work should develop multi-objective scaling frameworks that balance these factors to guide architecture and dataset design more holistically.
Inference-aware and modular scaling laws.
Traditional scaling laws assume fixed inference procedures. However, our synthesis in RQ4 and RQ7 shows that test-time compute allocation via sampling, retrieval, or routing can drastically affect performance. Future scaling formulations should modularize inference and allow flexible compute allocation per task or query.
Data quality over quantity.
Towards accessible and sustainable AI.
Large models are inaccessible to many research groups. Downscaling informed by scaling laws ensures that smaller labs and edge deployments can still benefit from state-of-the-art performance. Ultimately, the future of neural scaling is not just bigger models, but better modeling choices at every scale.
@@ 6 Conclusion
This survey provided a comprehensive analysis of neural scaling laws, exploring their theoretical foundations, empirical findings, and practical implications. It synthesized insights across various modalities, including language, vision, multimodal, and reinforcement learning, to uncover common trends and deviations from traditional power-law scaling. While early research established predictable relationships between model size, dataset volume, and computational resources, more recent studies have shown that these relationships are not universally applicable. Sparse architectures, retrieval-augmented models, and domain-specific adaptations often exhibit distinct scaling behaviors, challenging the notion of uniform scalability. Furthermore, advancements in fine-tuning, data pruning, and efficient inference strategies have introduced new perspectives on compute-optimal scaling. Despite their significance, scaling laws remain an evolving area of research, requiring further refinement to address real-world deployment challenges and architectural innovations.
@@
B Detailed scaling laws
- B.1 Scaling laws of language models
- B.2 Scaling laws in other modalities
- B.3 Scaling laws for domain adaptation
- B.4 Scaling laws for model inference
- B.5 Scaling laws for efficient models
- B.6 Data scaling laws
- B.7 Scaling laws for reinforcement learning
- B.8 Scaling laws for sparse autoencoders
- B.9 Scaling laws for graph neural networks
- C Reproducibility of scaling laws papers
@@ A.5 Limitations of fitting techniques
Appendix B Detailed scaling laws
B.1 Scaling laws of language models
| Modality | Paper | Key insights | Applicability |
[…]
B.2 Scaling laws in other modalities
| Paper | Key insights | Applicability |
| Zhai et al. (2022) | ViTs follow power-law scaling but plateau at extreme compute levels, with benefits primarily seen in datasets >1B images. | Image classification, object detection, large-scale vision datasets. |
| Aghajanyan et al. (2023) | Multimodal models experience competition at smaller scales but transition into synergy as model and token count grow, following a "competition barrier." | Multimodal learning, mixed-modal generative models, cross-domain AI. |
| Li et al. (2024a) | Scaling vision encoders in vision-language models (VLMs) does not always improve performance, reinforcing the importance of data quality over raw scaling. | Vision-language models, image-text alignment, multimodal scaling challenges. |
B.3 Scaling laws for domain adaptation
Fine-tuning strategies scale differently depending on dataset size. Parameter-efficient fine-tuning (PEFT) techniques like low-rank adaptation (LoRA) (Hu et al., 2021) and Prompt-tuning, both are well-suited for small datasets, but LoRA performs best for mid-sized datasets, and full fine-tuning is most effective for large datasets. However, PEFT methods provide better generalization in large models, making them attractive alternatives to full-scale fine-tuning (Zhang et al., 2024).
| Paper | Key insights | Applicability |
[…]
B.4 Scaling laws for model inference
| Paper | Key insights | Applicability |
[…]
B.5 Scaling laws for efficient models
| Paper | Key insights | Applicability |
| Clark et al. (2022) | Routing-based models show diminishing returns at larger scales, requiring optimal routing strategies for efficiency. | Routing-based models, MoEs, transformer scaling. |
| Krajewski et al. (2024) | Fine-grained MoEs achieve up to 40 compute efficiency gains when expert granularity is optimized. | Mixture of Experts models, large-scale compute efficiency. |
| Frantar et al. (2023) | Sparse model scaling enables predicting optimal sparsity levels for given compute budgets. | Sparse models, structured sparsity optimization, parameter reduction. |
B.6 Data scaling laws
| Paper | Key insights | Applicability |
[…]
B.7 Scaling laws for reinforcement learning
@@
B.8 Scaling laws for sparse autoencoders
B.9 Scaling laws for graph neural networks
Unlike in computer vision and natural language processing, where larger datasets typically improve generalization, graph self-supervised learning methods fail to exhibit expected scaling behavior and performance fluctuates unpredictably across different data scales (Ma et al., 2024). However, self-supervised learning pretraining loss does scale with more training data, but this improvement does not translate to better downstream performance. The scaling behavior is method-specific, with some approaches like InfoGraph showing more stable scaling than others like GraphCL.
Appendix C Reproducibility of scaling laws papers